Simplify the expression.
7P2/5P4

Answer by dabanfield(803)

7P2/5P4 =
(7/5)*(p2/p4) =
(7/5)*p^(2-4) =
(7/5)*(p^-2) =
(7/5)*(1/p^2) =
7/(5*p^2)